OpenStudy (saraa23):

How would you find the exact value? Cos(2arccos(5/7)) using double angle formula

OpenStudy (mathmale):

Sara: I don't see any "theta" in this problem statement. Please correct your statement and resubmit it here.

OpenStudy (saraa23):

I fixed the question. I added the theta on accident to the question

OpenStudy (anonymous):

\(\cos(2 \theta) = 2\cos^{2} (\theta) - 1\) Let your "theta" be equal to \(\arccos (5/7)\) and plug in.

OpenStudy (saraa23):

I am getting -48/49, and it is coming out to be wrong. I don't know where I'm going wrong

OpenStudy (danjs):

2 * [ cos( cos^-1( 5/7 )) ] ^2 - 1

OpenStudy (danjs):

that is 1/49 if that works

OpenStudy (danjs):

cos[ cos^-1(x)] ---> x , a function and inverse remove each other

OpenStudy (danjs):

2*[5/7]^2 - 1 2* (25/49) - 1 1/49

OpenStudy (saraa23):

Okay now I see why cos and cos^1 cancel out ! Makes so much more sense thank you !

OpenStudy (danjs):

i think you figured (5/7)^2 - 1, then doubled that, operation order little off

OpenStudy (danjs):

need 2*[5/7]^2 first , then take the 1 off

OpenStudy (saraa23):

Yess that's what I did. I see my mistake now.

OpenStudy (danjs):

k, welcome